OpenStudy (anonymous):

The sum of two numbers in 30 and their difference is 12. Write two equations and find the numbers.

OpenStudy (ajprincess):

Let us assume the numbers are x and y. Sum of them is the addition of then. so Sum = x+y It is given that the sum is equal to 30. so x+y=30 difference is subtraction of a number from the other. so difference=x-y it is given that the difference is equal to 12. so x-y=12. so the required equations are x+y=30 and x-y=12. getting it @oly.surf

OpenStudy (anonymous):

yep ive got that!

OpenStudy (anonymous):

x+y=30 x-y=12 on solving.. x=21 and y=9

OpenStudy (anonymous):

how did you solve that @chronics

OpenStudy (anonymous):

just add the two equations..y would cancel out and u get 2x=42 ..hence x=21..

OpenStudy (anonymous):

thanks